# Relativistic Roche problem for stars in precessing orbits around a spinning black hole

### v2

@@ I Introduction
In this paper, we study the case of a star in a circular orbit on the non-equatorial plane around a spinning BH. To study this, we assume that the star is incompressible and utilize the tensor virial formalism developed by Chandrasekhar [2]. This formalism has generally been used to study the assumed ellipsoidal shape of a self-gravitating fluid and provided a number of knowledge on the equilibrium and stability of the stars.
After reviewing the previous results, we show the presence of a resonance when the orbit has an inclination with respect to the equatorial plane. To compute the impact of the inclination on the shape of the star, we use a perturbative expansion assuming that the inclination angle is small: We perform this expansion up to the second order in the angle of the inclined orbit $\varepsilon=\theta-\pi/2$ and show that the resulting equations take a linear form, which is usually solvable. When this is not the case, we posit a new kind of resonance for such systems. We show that the resonances can play an important role in the evolution of BH-neutron star (NS) binaries in close orbits or of a white dwarf/an ordinary star orbiting a supermassive BH.
This paper is organized as follows: In Sec. II, we present the working assumptions and the tensor virial equations adapted to our system, by which we derive the equilibrium solution of a star on an equatorial orbit around a BH. In Sec. III, we present our method for the perturbative expansion, which we apply up to the second order in $\varepsilon$. In Sec. IV, we show the presence of new resonances in these systems, using the equations derived in Sec. III. Finally, we discuss the implications of our findings in Sec. V. Throughout this paper, we use the geometrical units of $c=G=1$ where $c$ and $G$ are the speed of light and gravitational constant, respectively. We will also not use the Einstein summation convention.
@@ II.1 System
The purpose of this paper is to study the tidal effect on a star orbiting a Kerr BH. We assume that the star is modeled by an incompressible fluid of mass $M$ and constant density $\rho$. This star is assumed to have an ellipsoidal shape determined by its principal semi-axes $(a_{1},a_{2},a_{3})$. We assume that it orbits a spinning BH of mass $m$ and of spin parameter $a$. When referencing its position relative to the BH, we will use the Boyder-Lindquist coordinates in which the line element is written as
@@
The expression of the tidal tensor can be simplified by changing the frame we are working on. Indeed in a rotating frame of rotation $\vec{\Omega}$ along the 3-axis with magnitude $\Omega=d\Psi/d\tau$, the tidal tensor becomes [10]
@@
$+6ar\cos\theta\frac{ST}{K\Sigma^{2}}I_{b},$
@@
$-6ar\cos\theta\frac{ST}{K\Sigma^{2}}I_{b},$
@@
$+(a^{2}\cos^{2}\theta S-r^{2}T)I_{b}\Big{]}\frac{\sqrt{ST}}{K\Sigma^{2}},$
@@ II.2 Tensor virial equations
As already stated in Sec. I, we employ the tensor virial equations [16, 2] to study the equilibrium state of stars orbiting a Kerr BH. The hydrodynamic equation for the fluid on the local inertial frame is written as
@@
where $X_{i}=(X_{1},X_{2},X_{3})$ denote the coordinates in the local inertial frame, $U_{i}$ is the velocity field of the star, $P$ the pressure, and $\phi$ the Newtonian potential of the fluid. It obeys the Poisson equation as
@@
As seen above, it is preferable to place ourselves in a rotating frame of rotation $\vec{\Omega}$, as this simplifies the expression of the tidal tensor. In this case, we have
@@
where $p_{c}$ is the central pressure and $a_{i}$ ($i=1$–3) denote the axial length of the star for the corresponding directions. This setting enables us to assume that the velocity depends linearly on the position as
@@
We now follow the method developed by Chandrasekhar to derive the second-rank tensor virial tensor equations [2]. First, we multiply $x_{j}$ to Eq. (23) and integrate over the volume $V$ of the entire star. Then we obtain
@@ II.3 Equatorial plane case
In this paper, we pay attention to the cases of $f=0$ and $f=1$. For $f=0$, the fluid has no internal motion in the rotating frame. In the inertial frame, the fluid’s internal motion comes only from the rotation. On the other hand, for $f=1$, the fluid has a certain vorticity in the rotating frame, while in the inertial frame, the fluid does not have any circulation. We refer to this case as irrotational Roche-Riemann ellipsoid (IRRE) [2]. It has been shown that this velocity configuration is typically (approximately) satisfied for binary neutron stars and BH-NS binaries in the late inspiraling stage [17, 18].
@@
One notices that $\tilde{\Omega}$ reaches a maximum value $\tilde{\Omega}_{\mathrm{crit}}$ corresponding to $\alpha_{3,\mathrm{crit}}$. Above this value, there is no possible equilibrium configuration of the star. This gives the relativistic Roche limit, already computed in [5, 8]. As in [8], we find that at the innermost stable circular orbit (ISCO) $\tilde{\Omega}_{\mathrm{crit}}^{2}\approx 0.06364$ for $f=1$, and $\tilde{\Omega}_{\mathrm{crit}}^{2}\approx 0.06640$ for $f=0$ [5]. As in the classical case, one can expect that the star is unstable against mass shedding/tidal disruption for $\alpha_{3}<\alpha_{3,\mathrm{crit}}$ and stable when $\alpha_{3}>\alpha_{3,\mathrm{crit}}$. Thus, for a given value of $\rho$, $\Omega$ has a maximum value, i.e., the orbital separation $r_{0}$ has a minimum value, and for the smaller orbital separation, the star should start the tidal disruption process.
@@ III.1 Method for the expansion
The displacement will depend on time, as the orbital value of $\theta$ does too. Thus, as a first step, we have to determine the time-dependent orbital motion. To do this we analyze the geodesic equation for $\theta$, (4), which is rewritten for $\varepsilon=\theta-\pi/2\ll 1$ as
@@
We consider a spherical orbit with always the same radius $r_{0}$ but different values of $\theta$. Following [14], to study the orbit of the star, we fix $r_{0}$ and $\cal{C}$ and deduce the other parameters from ${\cal{R}}=0=d{\cal{R}}/dr$ at $r=r_{0}$. Thus, $E$, $L$, and $K$ depend on the inclination angle $\theta$. These changes occur at the second order in $\varepsilon_{0}$ (see appendix).
@@
Following [2], the integrals following a displacement $\xi_{i}$ of the fluid are given as
@@
The resulting equations can be divided in two types as shown in the following subsections. In the following, we only consider equations which contain non-zero contributions of lower-order solutions of the fluid. This allows us to focus only on the displacement induced by the inclination of the orbit.
@@ III.2 First-order calculations
$\begin{split}&(-\omega_{\theta}^{2}+4\pi\rho B_{13}-\Omega_{0}^{2}+\tilde{C}_{11}+\tilde{C}_{33})V_{1;3}^{(1)}\\ &+(-\omega_{\theta}^{2}+4\pi\rho B_{13}-\Omega_{0}^{2}-2\Lambda^{2}\\ &+2\Omega_{0}\Lambda\alpha_{2}+\tilde{C}_{11}+\tilde{C}_{33})V_{3;1}^{(1)}\\ &-2\Omega_{0}\omega_{\theta}V_{2;3}^{(1)}-2\Lambda\omega_{\theta}\alpha_{2}^{-1}V_{3;2}^{(1)}=-M\tilde{C}_{13}^{(1)}\frac{a_{3}^{2}+a_{1}^{2}}{5},\end{split}$ (69)
$\begin{split}&(-\omega_{\theta}^{2}+4\pi\rho B_{23}-\Omega_{0}^{2}+\tilde{C}_{22}+\tilde{C}_{33})V_{2;3}^{(1)}\\ &+(-\omega_{\theta}^{2}+4\pi\rho B_{23}-\Omega_{0}^{2}-2\Lambda^{2}\\ &+2\Omega_{0}\Lambda\alpha_{2}^{-1}+\tilde{C}_{22}+\tilde{C}_{33})V_{3;2}^{(1)}\\ &-2\Omega_{0}\omega_{\theta}V_{1;3}^{(1)}-2\Lambda\omega_{\theta}\alpha_{2}V_{3;1}^{(1)}=0,\end{split}$ (70)
@@
$\begin{split}&(-\omega_{\theta}^{2}-\Omega_{0}^{2}+\tilde{C}_{11}-\tilde{C}_{33})V_{1;3}^{(1)}\\ &+(\omega_{\theta}^{2}-\Omega_{0}^{2}+2\Omega_{0}\Lambda\alpha_{2}+\tilde{C}_{11}-\tilde{C}_{33})V_{3;1}^{(1)}\\ &-2\Omega_{0}\omega_{\theta}V_{2;3}^{(1)}+2\Lambda\omega_{\theta}\alpha_{2}^{-1}V_{3;2}^{(1)}=-M\tilde{C}_{13}^{(1)}\frac{a_{3}^{2}-a_{1}^{2}}{5},\end{split}$ (71)
$\begin{split}&(-\omega_{\theta}^{2}-\Omega_{0}^{2}+\tilde{C}_{22}-\tilde{C}_{33})V_{2;3}^{(1)}\\ &+(\omega_{\theta}^{2}-\Omega_{0}^{2}+2\Omega_{0}\Lambda\alpha_{2}^{-1}+\tilde{C}_{22}-\tilde{C}_{33})V_{3;2}^{(1)}\\ &-2\Omega_{0}\omega_{\theta}V_{1;3}^{(1)}+2\Lambda\omega_{\theta}\alpha_{2}V_{3;1}^{(1)}=0.\end{split}$ (72)
@@ III.3 Second-order calculations
To find the second-order equations, we need to take into account second-order corrections to $\Omega$ and $\dot{\Omega}$ for the equations. We denote the former as $\Omega=\Omega_{0}+\Omega^{(2)}+O(\varepsilon_{0}^{4})$. The second-order corrections to $\Omega$ and $\dot{\Omega}$ are given in the appendix A.
@@
$=-\tilde{C}_{11}^{(2)}I_{11}+\tilde{C}_{33}^{(2)}I_{33}+4\Omega^{(2)}(\Omega_{0}-\Lambda\alpha_{2})I_{11},$ (82)
$\begin{split}&(-2\pi\rho(B_{13}-B_{12})+2\Omega_{0}\Lambda\alpha_{2})V_{1;1}\\ &+\Big{(}\frac{d^{2}}{d\tau^{2}}+2\pi\rho(3B_{22}-B_{23})-2\Omega_{0}^{2}-2\Lambda^{2}+2\Omega_{0}\Lambda\alpha_{2}^{-1}\\ &+2\tilde{C}_{22}\Big{)}V_{2;2}+\Big{(}-\frac{d^{2}}{d\tau^{2}}-2\pi\rho(3B_{33}-B_{23})-2\tilde{C}_{33}\Big{)}V_{3;3}\\ &+2\Omega_{0}\frac{d}{d\tau}V_{1;2}+2\Lambda\alpha_{2}\frac{d}{d\tau}V_{2;1}\\ &=-\tilde{C}_{22}^{(2)}I_{22}+\tilde{C}_{33}^{(2)}I_{33}+\tilde{C}_{13}(V_{3;1}+V_{1;3})\\ &+4\Omega^{(2)}(\Omega_{0}-\Lambda\alpha_{2}^{-1})I_{22},\end{split}$ (83)
$\begin{split}&(2\Lambda\alpha_{2}+2\Omega_{0})\frac{d}{d\tau}V_{1;1}-(2\Omega_{0}+2\Lambda\alpha_{2}^{-1})\frac{d}{d\tau}V_{2;2}\\ &+\Big{(}\frac{d^{2}}{d\tau^{2}}+4\pi\rho B_{12}-2\Omega_{0}^{2}-2\Lambda^{2}+\tilde{C}_{11}+\tilde{C}_{22}\Big{)}V_{1;2}\\ &+\Big{(}\frac{d^{2}}{d\tau^{2}}+4\pi\rho B_{12}-2\Omega_{0}^{2}-2\Lambda^{2}+\tilde{C}_{11}+\tilde{C}_{22}\Big{)}V_{2;1}\\ &=-\tilde{C}_{13}(V_{3;2}+V_{2;3})+\dot{\Omega}(I_{22}-I_{11}),\end{split}$ (84)
$\begin{split}&(2\Lambda\alpha_{2}-2\Omega_{0})\frac{d}{d\tau}V_{1;1}-(2\Omega_{0}-2\Lambda\alpha_{2}^{-1})\frac{d}{d\tau}V_{2;2}\\ &+\Big{(}\frac{d^{2}}{d\tau^{2}}+\tilde{C}_{11}-\tilde{C}_{22}\Big{)}V_{1;2}\\ &+\Big{(}-\frac{d^{2}}{d\tau^{2}}+\tilde{C}_{11}-\tilde{C}_{22}\Big{)}V_{2;1}\\ &=-\tilde{C}_{13}(V_{3;2}+V_{2;3})+\dot{\Omega}(I_{22}+I_{11}),\end{split}$ (85)
$V_{1;1}+\frac{1}{\alpha_{2}^{2}}V_{2;2}+\frac{1}{\alpha_{3}^{2}}V_{3;3}=0.$ (86)
@@
$\begin{split}\tilde{C}_{11}^{(2)}&=\tilde{C}_{11}^{(2,0)}+\tilde{C}_{11}^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \tilde{C}_{22}^{(2)}&=\tilde{C}_{22}^{(2,0)}+\tilde{C}_{22}^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \tilde{C}_{33}^{(2)}&=\tilde{C}_{33}^{(2,0)}+\tilde{C}_{33}^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \Omega^{(2)}&=\Omega^{(2,0)}+\Omega^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \dot{\Omega}&=\dot{\Omega}^{(2)}\sin(2\omega_{\theta}\tau),\\ \tilde{C}_{13}(V_{3;1}+V_{1;3})&=\frac{\tilde{C}_{13}^{(1)}(V_{3;1}^{(1)}+V_{1;3}^{(1)})}{2}[1+\cos(2\omega_{\theta}\tau)],\\ \tilde{C}_{13}(V_{3;2}+V_{2;3})&=\frac{\tilde{C}_{13}^{(1)}(V_{3;2}^{(1)}+V_{2;3}^{(1)})}{2}\sin(2\omega_{\theta}\tau).\\ \end{split}$ (88)
@@
$\begin{split}C^{(2,2)}_{1}=&-\tilde{C}_{11}^{(2,2)}I_{11}+\tilde{C}_{33}^{(2,2)}I_{33}+4\Omega^{(2,2)}(\Omega_{0}-\Lambda\alpha_{2})I_{11},\\ C^{(2,2)}_{2}=&-\tilde{C}_{22}^{(2,2)}I_{22}+\tilde{C}_{33}^{(2,2)}I_{33}+\tilde{C}_{13}^{(1)}(V_{3;1}^{(1)}+V_{1;3}^{(1)})\\ &+4\Omega^{(2,2)}(\Omega_{0}-\Lambda\alpha_{2}^{-1})I_{22},\\ C^{(2,2)}_{3}=&-\frac{\tilde{C}_{13}^{(1)}(V_{3;2}^{(1)}+V_{2;3}^{(1)})}{2}+\dot{\Omega}^{(2)}(I_{22}-I_{11}),\\ C^{(2,2)}_{4}=&-\frac{\tilde{C}_{13}^{(1)}(V_{3;2}^{(1)}+V_{2;3}^{(1)})}{2}+\dot{\Omega}(I_{22}+I_{11}),\\ C^{(2,2)}_{5}=&0.\\ \end{split}$ (95)
The differential equations for the constant term must be considered with more care. If we proceed as before, we obtain again linear equations $\sum_{k=1}^{3}M_{ik}^{(2,0)}V_{k}^{(2,0)}=C_{i}^{(2,0)}$. The matrix of this equation $M_{ik}^{(2,0)}$ is the same as $M_{ik}^{(2,2)}$ when we set $\omega_{\theta}=0$. However, this matrix is never invertible as its third and fourth line are colinear.

### v3

@@ I Introduction
In this paper, we study the case of a star in a circular orbit on the non-equatorial plane around a spinning BH. To study this, as a first step, we assume that the star is incompressible and utilize the tensor virial formalism developed by Chandrasekhar [2]. This formalism has generally been used to study the assumed ellipsoidal shape of a self-gravitating fluid and provided a number of knowledge on the equilibrium and stability of the stars qualitatively.
After reviewing the previous results, we show the presence of a resonance when the orbit has an inclination with respect to the equatorial plane. To compute the impact of the inclination on the shape of the star, we use a perturbative expansion assuming that the inclination angle is small: We perform this expansion up to the second order in the angle of the inclined orbit $\varepsilon=\theta-\pi/2$ and show that the resulting equations take a linear form, which is usually solvable. When this is not the case, we posit a new kind of a resonance for such systems. We show that the resonances can play an important role in the evolution of BH-neutron star (NS) binaries in close orbits or of a white dwarf/an ordinary star orbiting a supermassive BH.
This paper is organized as follows: In Sec. II, we present the working assumptions and the tensor virial equations adapted to our system, by which we derive the equilibrium solution of a star on an equatorial orbit around a BH [5, 6, 7, 8]. In Sec. III, we present our method for the perturbative expansion, which we apply up to the second order in $\varepsilon$. In Sec. IV, we show the presence of new resonances in these systems, using the equations derived in Sec. III. Finally, we discuss the implications of our findings in Sec. V. Throughout this paper, we use the geometrical units of $c=G=1$ where $c$ and $G$ are the speed of light and gravitational constant, respectively. We will not use the Einstein summation convention.
@@ II.1 System
The purpose of this paper is to study the tidal effect on a star orbiting a Kerr BH. We assume that the star is modeled by an incompressible fluid of mass $M$ and constant density $\rho$. This star is assumed to have an ellipsoidal shape determined by its principal semi-axes $(a_{1},a_{2},a_{3})$. We assume that it orbits a spinning BH of mass $m$ and of spin parameter $a$. When referencing its position relative to the BH, we will use the Boyer-Lindquist coordinates in which the line element is written as
@@
The expression of the tidal tensor can be simplified by changing the frame we are working on. Indeed in a rotating frame of rotation $\vec{\Omega}$ along the 3-axis with magnitude $\Omega=d\Psi/d\tau$, the tidal tensor is simplified as [10]
@@
$~{}+6ar\cos\theta\frac{ST}{K\Sigma^{2}}I_{b},$
@@
$~{}-6ar\cos\theta\frac{ST}{K\Sigma^{2}}I_{b},$
@@
$~{}+(a^{2}\cos^{2}\theta S-r^{2}T)I_{b}\Big{]}\frac{\sqrt{ST}}{K\Sigma^{2}},$
@@ II.2 Tensor virial equations
As already stated in Sec. I, we employ the tensor virial equations [16, 2] to study the equilibrium state of stars orbiting a Kerr BH. The hydrodynamics equation for the fluid on the local inertial frame is written as
@@
where $X_{i}=(X_{1},X_{2},X_{3})$ denote the coordinates in the local inertial frame, $U_{i}$ is the velocity field of the star, $P$ the pressure, and $\phi$ the Newtonian potential of the fluid, which obeys the Poisson equation as
@@
As seen above, it is preferable to work in a rotating frame of angular velocity vector $\vec{\Omega}$, as this simplifies the expression of the tidal tensor. In this case, the equation of motion is rewritten as
@@
where $p_{c}$ is the central pressure and $a_{i}$ ($i=1$–3) denote the axial lengths of the star for the corresponding directions. This setting enables us to assume that the velocity depends linearly on the position as
@@
We now follow the method developed by Chandrasekhar to derive the second-rank tensor virial equations [2]. First, we multiply $x_{j}$ to Eq. (23) and integrate over the volume $V$ of the entire star. Then we obtain
@@ II.3 Equatorial plane case
In this paper, we pay attention to the cases of $f=0$ and $f=1$. For $f=0$, the fluid has no internal motion in the rotating frame, while in the inertial frame, the fluid’s internal motion comes only from the rotation. On the other hand, for $f=1$, the fluid has a certain vorticity in the rotating frame, while in the inertial frame, the fluid does not have any circulation. We refer to this case as irrotational Roche-Riemann ellipsoid (IRRE) [2]. It has been shown that this velocity configuration is typically (approximately) satisfied for binary neutron stars and BH-NS binaries in the late inspiraling stage [17, 18].
@@
One notices that $\tilde{\Omega}$ reaches a maximum value $\tilde{\Omega}_{\mathrm{crit}}$ corresponding to $\alpha_{3,\mathrm{crit}}$. Above this value, there is no possible equilibrium configuration of the star. This gives the relativistic Roche limit, already computed in [5, 8]. As in [8], we find that at the innermost stable circular orbit (ISCO) $\tilde{\Omega}_{\mathrm{crit}}^{2}\approx 0.06364$ for $f=1$, and $\tilde{\Omega}_{\mathrm{crit}}^{2}\approx 0.06640$ for $f=0$ [5]. As in the classical case, one can expect that the star is unstable against mass shedding/tidal disruption for $\alpha_{3}<\alpha_{3,\mathrm{crit}}$ and stable for $\alpha_{3}>\alpha_{3,\mathrm{crit}}$. Thus, for a given value of $\rho$, $\Omega$ has a maximum value, i.e., the orbital separation $r_{0}$ has a minimum value, and for the smaller orbital separation, the star should start the tidal disruption process.
@@ III.1 Method for the expansion
By perturbing the equatorial solution, we do not provide any change to the Roche limit for stars in spherical orbit as in [9]. However this method enables us to study more in detail the dynamical behavior of the star. In particular we do not restrict the motion of the fluid to be in the 1-2 plane. While we assume such a motion in the equatorial case, we should expect a wider range of motions due to the non-diagonal elements of the tidal tensor. This intuition will be confirmed in the following analyses.
@@
The displacement will depend on time, as the orbital value of $\theta$ does too. Thus, as a first step, we have to determine the time-dependent orbital motion. To do this we analyze the geodesic equation for $\theta$, Eq. (4), which is rewritten for $\varepsilon=\theta-\pi/2\ll 1$ as (e.g., [19])
@@
We consider a spherical orbit with always the same radius $r_{0}$ but different values of $\theta$. Following [14], to study the orbit of the star, we fix $r_{0}$ and $\cal{C}$ and deduce the other parameters from ${\cal{R}}=0=d{\cal{R}}/dr$ at $r=r_{0}$. Thus, $E$, $L$, and $K$ depend on the inclination angle $\varepsilon_{0}$. These changes occur at the second order in $\varepsilon_{0}$ (see appendix).
@@
Following [2], the integrals associated with a displacement $\xi_{i}$ of the fluid are given as
@@
The resulting equations can be divided in two types as shown in the subsequent subsections. In the following, we only consider equations which contain non-zero contributions of lower-order solutions of the fluid. This allows us to focus only on the displacement induced by the inclination of the orbit.
@@ III.2 First-order calculations
$(-\omega_{\theta}^{2}+4\pi\rho B_{13}-\Omega_{0}^{2}+\tilde{C}_{11}+\tilde{C}_{33})V_{1;3}^{(1)}$
$+(-\omega_{\theta}^{2}+4\pi\rho B_{13}-\Omega_{0}^{2}-2\Lambda^{2}$
$+2\Omega_{0}\Lambda\alpha_{2}+\tilde{C}_{11}+\tilde{C}_{33})V_{3;1}^{(1)}$
$-2\Omega_{0}\omega_{\theta}V_{2;3}^{(1)}-2\Lambda\omega_{\theta}\alpha_{2}^{-1}V_{3;2}^{(1)}=-M\tilde{C}_{13}^{(1)}\frac{a_{3}^{2}+a_{1}^{2}}{5},~{}~{}$ (69)
$(-\omega_{\theta}^{2}+4\pi\rho B_{23}-\Omega_{0}^{2}+\tilde{C}_{22}+\tilde{C}_{33})V_{2;3}^{(1)}$
$+(-\omega_{\theta}^{2}+4\pi\rho B_{23}-\Omega_{0}^{2}-2\Lambda^{2}$
$+2\Omega_{0}\Lambda\alpha_{2}^{-1}+\tilde{C}_{22}+\tilde{C}_{33})V_{3;2}^{(1)}$
$-2\Omega_{0}\omega_{\theta}V_{1;3}^{(1)}-2\Lambda\omega_{\theta}\alpha_{2}V_{3;1}^{(1)}=0,$ (70)
@@
$(-\omega_{\theta}^{2}-\Omega_{0}^{2}+\tilde{C}_{11}-\tilde{C}_{33})V_{1;3}^{(1)}$
$+(\omega_{\theta}^{2}-\Omega_{0}^{2}+2\Omega_{0}\Lambda\alpha_{2}+\tilde{C}_{11}-\tilde{C}_{33})V_{3;1}^{(1)}$
$-2\Omega_{0}\omega_{\theta}V_{2;3}^{(1)}+2\Lambda\omega_{\theta}\alpha_{2}^{-1}V_{3;2}^{(1)}=-M\tilde{C}_{13}^{(1)}\frac{a_{3}^{2}-a_{1}^{2}}{5},$ (71)
$(-\omega_{\theta}^{2}-\Omega_{0}^{2}+\tilde{C}_{22}-\tilde{C}_{33})V_{2;3}^{(1)}$
$+(\omega_{\theta}^{2}-\Omega_{0}^{2}+2\Omega_{0}\Lambda\alpha_{2}^{-1}+\tilde{C}_{22}-\tilde{C}_{33})V_{3;2}^{(1)}$
$-2\Omega_{0}\omega_{\theta}V_{1;3}^{(1)}+2\Lambda\omega_{\theta}\alpha_{2}V_{3;1}^{(1)}=0.$ (72)
@@ III.3 Second-order calculations
To find the second-order equations, we need to take into account second-order corrections to $\Omega$ and $\dot{\Omega}$ for the equations. We denote the former as $\Omega=\Omega_{0}+\Omega^{(2)}+O(\varepsilon_{0}^{4})$. The second-order corrections to $\Omega$ and $\dot{\Omega}$ are given in the appendix B.
One must not forget quadratic contributions from the first order perturbations. Indeed the total moment of inertia tensor computed up to the second order is
$I^{tot}_{ij}$ $=$ $\int_{V}(x_{i}+\xi^{(1)}_{ik}x_{k}+\xi^{(2)}_{ik}x_{k})(x_{j}+\xi^{(1)}_{jl}x_{l}+\xi^{(2)}_{jl}x_{l})d^{3}x$ (81)
$=$ $I^{(0)}_{ij}+V^{(1)}_{ij}+V^{(2)}_{ij}+\frac{5}{Ma_{k}^{2}}V_{i;k}^{(1)}V^{(1)}_{j;k}+O(\varepsilon^{3}).$
We will in the following abbreviate these contributions to $Q_{ij}(V^{(1)}_{k;l})$, while giving them their complete form in the appendix B. Given our displacement, these terms are non-zero only when $(i,j)$ is even on the component 3.
@@
$=-\tilde{C}_{11}^{(2)}I_{11}+\tilde{C}_{33}^{(2)}I_{33}+4\Omega^{(2)}(\Omega_{0}-\Lambda\alpha_{2})I_{11}+Q_{11}(V^{(1)}_{k;l}),$
$(-2\pi\rho(B_{13}-B_{12})+2\Omega_{0}\Lambda\alpha_{2})V_{1;1}$
$+\Big{(}\frac{d^{2}}{d\tau^{2}}+2\pi\rho(3B_{22}-B_{23})-2\Omega_{0}^{2}-2\Lambda^{2}+2\Omega_{0}\Lambda\alpha_{2}^{-1}$
$+2\tilde{C}_{22}\Big{)}V_{2;2}+\Big{(}-\frac{d^{2}}{d\tau^{2}}-2\pi\rho(3B_{33}-B_{23})-2\tilde{C}_{33}\Big{)}V_{3;3}$
$+2\Omega_{0}\frac{d}{d\tau}V_{1;2}+2\Lambda\alpha_{2}\frac{d}{d\tau}V_{2;1}$
$=-\tilde{C}_{22}^{(2)}I_{22}+\tilde{C}_{33}^{(2)}I_{33}+\tilde{C}_{13}(V_{3;1}+V_{1;3})$
$~{}~{}+4\Omega^{(2)}(\Omega_{0}-\Lambda\alpha_{2}^{-1})I_{22}+Q_{22}(V^{(1)}_{k;l}),$ (84)
$(2\Lambda\alpha_{2}+2\Omega_{0})\frac{d}{d\tau}V_{1;1}-(2\Omega_{0}+2\Lambda\alpha_{2}^{-1})\frac{d}{d\tau}V_{2;2}$
$+\Big{(}\frac{d^{2}}{d\tau^{2}}+4\pi\rho B_{12}-2\Omega_{0}^{2}-2\Lambda^{2}+\tilde{C}_{11}+\tilde{C}_{22}\Big{)}V_{1;2}$
$+\Big{(}\frac{d^{2}}{d\tau^{2}}+4\pi\rho B_{12}-2\Omega_{0}^{2}-2\Lambda^{2}+\tilde{C}_{11}+\tilde{C}_{22}\Big{)}V_{2;1}$
$=-2\tilde{C}_{13}(V_{3;2}+V_{2;3})+\dot{\Omega}(I_{22}-I_{11})$
$+Q_{12}(V^{(1)}_{k;l})+Q_{21}(V^{(1)}_{k;l}),$ (85)
$(2\Lambda\alpha_{2}-2\Omega_{0})\frac{d}{d\tau}V_{1;1}-(2\Omega_{0}-2\Lambda\alpha_{2}^{-1})\frac{d}{d\tau}V_{2;2}$
$+\Big{(}\frac{d^{2}}{d\tau^{2}}+\tilde{C}_{11}-\tilde{C}_{22}\Big{)}V_{1;2}$
$+\Big{(}-\frac{d^{2}}{d\tau^{2}}+\tilde{C}_{11}-\tilde{C}_{22}\Big{)}V_{2;1}$
$=\dot{\Omega}(I_{22}+I_{11})+Q_{12}(V^{(1)}_{k;l})-Q_{21}(V^{(1)}_{k;l}),$ (86)
and
$V_{1;1}+\frac{1}{\alpha_{2}^{2}}V_{2;2}+\frac{1}{\alpha_{3}^{2}}V_{3;3}=\sum_{k=1}^{3}\sum_{l=1}^{3}\frac{V^{(1)}_{k;l}V^{(1)}_{l;k}}{2\alpha_{k}^{2}\alpha_{l}^{2}}.$ (87)
@@
$\begin{split}\tilde{C}_{11}^{(2)}&=\tilde{C}_{11}^{(2,0)}+\tilde{C}_{11}^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \tilde{C}_{22}^{(2)}&=\tilde{C}_{22}^{(2,0)}+\tilde{C}_{22}^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \tilde{C}_{33}^{(2)}&=\tilde{C}_{33}^{(2,0)}+\tilde{C}_{33}^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \Omega^{(2)}&=\Omega^{(2,0)}+\Omega^{(2,2)}\cos(2\omega_{\theta}\tau),\\ \dot{\Omega}&=\dot{\Omega}^{(2)}\sin(2\omega_{\theta}\tau),\\ \tilde{C}_{13}(V_{3;1}+V_{1;3})&=\frac{\tilde{C}_{13}^{(1)}(V_{3;1}^{(1)}+V_{1;3}^{(1)})}{2}[1+\cos(2\omega_{\theta}\tau)],\\ \tilde{C}_{13}(V_{3;2}+V_{2;3})&=\frac{\tilde{C}_{13}^{(1)}(V_{3;2}^{(1)}+V_{2;3}^{(1)})}{2}\sin(2\omega_{\theta}\tau).\end{split}$ (89)
@@
$\begin{split}C^{(2,2)}_{1}=&-\tilde{C}_{11}^{(2,2)}I_{11}+\tilde{C}_{33}^{(2,2)}I_{33}\\ &+4\Omega^{(2,2)}(\Omega_{0}-\Lambda\alpha_{2})I_{11}+Q_{11}^{(2,2)}(V^{(1)}_{k;l}),\\ C^{(2,2)}_{2}=&-\tilde{C}_{22}^{(2,2)}I_{22}+\tilde{C}_{33}^{(2,2)}I_{33}+\frac{\tilde{C}_{13}^{(1)}(V_{3;1}^{(1)}+V_{1;3}^{(1)})}{2}\\ &+4\Omega^{(2,2)}(\Omega_{0}-\Lambda\alpha_{2}^{-1})I_{22}+Q_{22}^{(2,2)}(V^{(1)}_{k;l}),\\ C^{(2,2)}_{3}=&-\tilde{C}_{13}^{(1)}(V_{3;2}^{(1)}+V_{2;3}^{(1)})+\dot{\Omega}^{(2)}(I_{22}-I_{11})\\ &+Q_{12}^{(2,2)}(V^{(1)}_{k;l})+Q_{21}^{(2,2)}(V^{(1)}_{k;l}),\\ C^{(2,2)}_{4}=&\dot{\Omega}(I_{22}+I_{11})+Q_{12}^{(2,2)}(V^{(1)}_{k;l})-Q_{21}^{(2,2)}(V^{(1)}_{k;l}),\\ C^{(2,2)}_{5}=&\sum_{k=1}^{3}\sum_{l=1}^{3}\frac{V^{(1)}_{k;l}V^{(1)}_{l;k}}{2\alpha_{k}^{2}\alpha_{l}^{2}}.\\ \end{split}$ (96)
The differential equations for the constant term must be considered with more care. If we proceed as before, we obtain again linear equations $\sum_{k=1}^{5}M_{ik}^{(2,0)}V_{k}^{(2,0)}=C_{i}^{(2,0)}$. The matrix of this equation $M_{ik}^{(2,0)}$ is the same as $M_{ik}^{(2,2)}$ when we set $\omega_{\theta}=0$. However, this matrix is never invertible as its third and fourth line are colinear.